Aggregator

CIS Work Summary 01/28/2026

6 days 19 hours ago
Log Number: 4467614Logbook(s): Body: 

- Did beam QE scan - max QE 1.36%
- Connected new HV Driver to RTP Pockets Cell
- PQB beam study transmission test for Aperture A2 at 4 mm was unsuccessful - will fix issues today.

suleiman

1L18 has been checked by Ops

6 days 19 hours ago
Log Number: 4467613Logbook(s): Entry Makers: msmithBody: 

Cav-1: Bypassed, not touching
Cav-2: steady at 4MV/m GSET
Cav-3: GSET 0, not touching
Cav-4: GSET 0, not touching
Cav-5: Steady at 3MV/m GSET after OLT
Cav-6: Bypassed, not touching
Cav-7: Steady at 4MV/m GSET
Cav-8: Steady at 4MV/m GSET

msmith

1L15 checked by Ops

6 days 20 hours ago
Log Number: 4467612Logbook(s): Entry Makers: msmithBody: 

All cavities recovered to 4 MV/m GSET.

Cav-2 recovered at first, then had ARC faults. Good after OLT.
Cav-3 required OLT.
Cav-6 struggling with phase and gradient loops. OLT did not help. Otherwise, stable at 4 MV/m.

msmith

booster cavity tuning and in SELAP

1 week ago
Log Number: 4467610Logbook(s): Body: 

cavities were pretty far off from the machine frequency
cavity 1 - ~24 kHz
cavity 2 - ~15 kHz

tuned the cavities closer to machine frequency. increased the gradients in SEL to the set points from the last run
cavity 1 - 2.5 MV/m
cavity 2 - 10.09 MV/m

made sure drive level (GASK) and the forward power levels (CRFP) are similar to the ones from the last run

poff optimization was done after tuning the cavities. used the one button to recover the cavities into SELAP

bachiman

Shift Information

Shift Owl
CEBAF Program
MCC Crew Chief Deir
MCC Operators Howard
Program Deputy Mock
Announcements

Access Status

Recent PSS History

recent pss history graph